# Emergence Learning: A Rising Direction from Emergent Abilities and a Monosemanticity-Based Study

### 17 Dec 2023 v1

@@ Appendix A Examples of Decreasing Monosemanticity in Large Models
$\theta_{u}\leftarrow\theta-l\nabla\mathcal{L}(\theta)=\theta+l\nabla f(\delta)\nabla g(\theta)$ (27)
By updating the parameter, we can express the updated $h$ with a variation on the multivariate Taylor expansion:
$h_{u}=$ $g(\theta_{u})$ (28)
$=$ $g(\theta)+l\nabla g(\theta)^{T}\nabla f(\delta)\nabla g(\theta)+l^{2}\frac{1}{2}(\nabla\mathcal{L}(\theta))^{T}\nabla^{2}g(\theta^{\prime})(\nabla\mathcal{L}(\theta)),$ (29)
where $\theta_{u}$ is the updated parameter and $\theta^{\prime}$ is between $\theta$ and $\theta_{u}$.
@@
$h_{u}-h_{o}=$ $l\nabla g(\theta)^{T}\nabla f(\delta)\nabla g(\theta)+\frac{l^{2}}{2}(\nabla\mathcal{L}(\theta))^{T}\nabla^{2}g(\theta^{\prime})(\nabla\mathcal{L}(\theta))$ (30)
$=$ $l(d_{1}+\frac{l}{2}d_{2}),$ (31)
which is greater than 0 when $d_{2}\geq 0,l>0$ or $d_{2}<0,0<l<\frac{2d_{1}}{-d_{2}}$, where $d_{1}=\nabla f(\delta)||\nabla g(\theta)||^{2}$ and $d_{2}=(\nabla\mathcal{L}(\theta))^{T}\nabla^{2}g(\theta^{\prime})(\nabla\mathcal{L}(\theta))$. As $t=h-\bar{h}<0$ under $\nabla f(\delta)\geq 0$, the activation value is smaller than the mean $\bar{h}$. Note that after the update, $h_{u}>h_{o}$. With a small enough learning rate $l$ (i.e., $<\frac{2d_{1}}{-d_{2}}$), $h_{u}$ will be updated towards $\bar{h}$ and deactivated (i.e., $\bar{h}>h_{u}>h_{o}$)
@@
$h_{u}-h_{o}=$ $l\nabla g(\theta)^{T}\nabla f(\delta)\nabla g(\theta)+\frac{l^{2}}{2}(\nabla\mathcal{L}(\theta))^{T}\nabla^{2}g(\theta^{\prime})(\nabla\mathcal{L}(\theta))$ (32)
$\leq$ $l\nabla f(\delta)||g(\theta)||^{2}+\frac{l^{2}L}{2}||\nabla g(\theta)||^{2},$ (33)
$=$ $l(\nabla f(\delta)+l\frac{L}{2})||\nabla g(\theta)||^{2}$ (34)
where Equation. (34) is valid as $g$ is the $h$-related part of model $\mathcal{F}$ and has a Lipschitz continuous gradient. Equation. (34) is smaller than 0 when learning rate $0<l<(-\nabla f(\delta))\frac{2}{L}$, which ensures that updated output $h_{u}$ is smaller than original activated value $h_{o}$, and thus close to $\bar{h}$ as $t=h-\bar{h}\geq 0$. Together with the case when $\nabla f(\delta)\geq 0$, we show that $h$ is always pushed to mean $\bar{h}$ and deactivated.
We point out that the loss function $\mathcal{L}(Z)$ does not always strictly increase with the deviation from $h$ (i.e., $|h^{\prime}-h|$) in neural networks. This is due to the presence of local optima, where increasing deviation from $h$ may actually lead to a smaller loss. Fortunately, in the case of a monosemantic neuron, the output is strongly influenced by $h$ and exhibits less nonlinearity, aligning more closely with the assumption.
Appendix D Implementation Details of MS $\phi$
@@
  Input: new batch of values $\boldsymbol{h}=\left\{h_{(m+1)},h_{(m+2)},\cdots,h_{(m+b)}\right\}$ of the neuron.
@@
  Calculate the MS for input values $\phi(h_{(m+i)})=(h_{(m+i)}-\mu_{m})^{2}/\upsilon_{m}$ for $i\in[1,b]$.
  Calculate $\mu_{b}^{\prime}=\frac{1}{b}\sum_{i=1}^{b}h_{(m+i)}$ and $\upsilon^{\prime}_{b}=\frac{1}{b}\sum_{i=1}^{b}(h_{(m+i)}-\mu^{\prime}_{b})^{2}$.
@@
  RETURN: Monosementicity Scale of inputs $\phi({\boldsymbol{h}})$
Appendix E Implementation Details of MEmeL
@@
  Input: Values of $n$ neurons with batchsize $b$: $H=\{\boldsymbol{h_{1}},\boldsymbol{h_{2}},\cdots,\boldsymbol{h_{n}}\}$, where $\boldsymbol{h_{i}}=\left\{{h_{i}}_{(m+1)},{h_{i}}_{(m+2)},\cdots,{h_{i}}_{(m+b)}\right\}$.
@@
$\phi({h_{i}}_{(m+j)})=\frac{({h_{i}}_{(m+j)}-{\mu_{i}}_{m})^{2}}{{\upsilon_{i}}_{m}+\epsilon\sum_{k=1}^{n}{\upsilon_{k}}_{m}/n}\text{ for }i\in[1,n],j\in[1,b].$
     Select values with high $\phi$, adjust each of them according to MD. Replacing the original ones in $H$ to form output $H^{\prime}$
@@
  RETURN: Adjusted layer $H^{\prime}$
Appendix F Experiment Setting
F.1 Benchmarks
@@
For the precipitation task, we utilize radar images with a granularity of $120\times 120$ from the HKO-7 dataset (Shi et al., 2017), which are centered in Hong Kong. These radar images cover a timespan from 2009 to 2015 and are captured every 6 minutes. The original rainfall intensity is mapped to a scale of [0, 255], forming a regression task. We will release the checkpoints and models soon  [^1] Codes are published and updating on https://github.com/dominatorX/MEmeL-code.
F.2 Metrics
@@
The Accuracy:
in a classification task, with the statistic of True Positive (TP), False Positive (FP), True Negative (TN), and False Negative (FN), Accuracy can be derived as $\frac{\text{TP+TN}}{\text{TP+TN+FN+FN}}$.
@@ Appendix G Discussions
G.1 Biological View of Neuron Connection
Monosemanticity and polysemanticity are inherent mappings between neurons. Similar to their functionality as connections, in neuroscience, “synaptic” enables information transmission between neurons (Hasson et al., 2020; Thomas et al., 2017). In view of the evolution of the brain, the increase of neuron activity is positively correlated to the increase of the amount of synaptic (M & S, 2015). Besides, compared with chimpanzees, humans significantly have more synaptics for information transmission (A et al., 2012).
However, the connections of artificial neurons are fixed during design. In contrast, the connection relationships of biological neurons are observable, which can serve as a reference for studying the monosemanticity of AI neurons in large-scale models.
G.2 Brain-inspired ANN design
@@
G.3 On the Selection of Benchmark and Experimental Settings
@@
However, validating the idea and performing pretraining is much more computationally expensive compared to finetuning, especially for large models. For instance, using the naive method to finetune LLAMA2-7B based on our A800 is estimated to cost over $450k$ and over 5800 GPU-days (Touvron et al., 2023) with our preliminary test. Emergence occurs primarily near $10^{22}$ FLOPs and costs more than 10500 GPU-days (Wei et al., 2022). Conducting corresponding validation experiments is impossible for most researchers without the assistance of top companies.
Under such an arms race on GPUs, our experiments explore our resources to validate the generality and effectiveness of our methods. For the three task domains, we perform fine-tuning on their milestone models using MEmeL, although this approach does not fully exploit the capabilities of MEmeL compared to pretraining.
G.4 Difference from Traditional Normalizations

### 15 Feb 2024 v2

@@ B.3. Debate on the Existence of Emergence
(33) $\theta^{*}\leftarrow\theta-l\nabla\mathcal{L}(\theta)=\theta+l\nabla f(\delta)\nabla g(\theta)$
By updating the parameter, we can express the updated $z$ with a variation on the multivariate Taylor expansion:
(34) $z^{*}=$ $g(\theta^{*})$
(35) $=$ $g(\theta)+l\nabla g(\theta)^{T}\nabla f(\delta)\nabla g(\theta)+l^{2}\frac{1}{2}(\nabla\mathcal{L}(\theta))^{T}\nabla^{2}g(\theta^{\prime})(\nabla\mathcal{L}(\theta)),$
where $\theta^{*}$ is the updated parameter and $\theta^{\prime}$ is between $\theta$ and $\theta^{*}$.
@@
(36) $z^{*}-z_{o}=$ $l\nabla g(\theta)^{T}\nabla f(\delta)\nabla g(\theta)+\frac{l^{2}}{2}(\nabla\mathcal{L}(\theta))^{T}\nabla^{2}g(\theta^{\prime})(\nabla\mathcal{L}(\theta))$
(37) $=$ $l(d_{1}+\frac{l}{2}d_{2}),$
which is greater than 0 when $d_{2}\geq 0,l>0$ or $d_{2}<0,0<l<\frac{2d_{1}}{-d_{2}}$, where $d_{1}=\nabla f(\delta)||\nabla g(\theta)||^{2}$ and $d_{2}=(\nabla\mathcal{L}(\theta))^{T}\nabla^{2}g(\theta^{\prime})(\nabla\mathcal{L}(\theta))$. As $t=z-\bar{z}<0$ under $\nabla f(\delta)\geq 0$, the activation value is smaller than the mean $\bar{h}$. Note that after the update, $z^{*}>z_{o}$. With a small enough learning rate $l$ (i.e., $<\frac{2d_{1}}{-d_{2}}$), $z^{*}$ will be updated towards $\bar{h}$ and deactivated (i.e., $\bar{h}>z^{*}>z_{o}$)
@@
(38) $z^{*}-z_{o}=$ $l\nabla g(\theta)^{T}\nabla f(\delta)\nabla g(\theta)+\frac{l^{2}}{2}(\nabla\mathcal{L}(\theta))^{T}\nabla^{2}g(\theta^{\prime})(\nabla\mathcal{L}(\theta))$
(39) $\leq$ $l\nabla f(\delta)||g(\theta)||^{2}+\frac{l^{2}L}{2}||\nabla g(\theta)||^{2},$
(40) $=$ $l(\nabla f(\delta)+l\frac{L}{2})||\nabla g(\theta)||^{2}$
where Equation. (40) is valid as $g$ is the $z$-related part of model $f$ and has a Lipschitz continuous gradient. Equation. (40) is smaller than 0 when learning rate $0<l<(-\nabla f(\delta))\frac{2}{L}$, which ensures that updated output $z^{*}$ is smaller than original activated value $z_{o}$, and thus close to $\bar{z}$ as $t=z-\bar{z}\geq 0$. Together with the case when $\nabla f(\delta)\geq 0$, we show that $z$ is always pushed to mean $\bar{z}$ and deactivated.
We point out that the loss function $\mathcal{L}(\mathbf{o})$ does not always strictly increase with the deviation from $z$ (i.e., $|z^{\prime}-z|$) in neural networks. This is due to the presence of local optima, where increasing deviation from $z$ may actually lead to a smaller loss. Fortunately, in the case of a monosemantic neuron, the output is strongly influenced by $z$ and exhibits less nonlinearity, aligning more closely with the assumption.
Appendix D Implementation Details
D.1. Implementation Details of MS $\phi$
@@
  Input: new batch of values $\left\{\mathbf{z}^{[m+1]},\mathbf{z}^{[m+2]},\cdots,\mathbf{z}^{[m+b]}\right\}$ of the neuron.
@@
  Calculate the MS for each input value $z\in\mathbf{z}$: $\phi(z^{[m+i]})=(z^{[m+i]}-\mu_{m})^{2}/\upsilon_{m}$ for $i\in[1,b]$.
  Calculate $\mu_{b}^{\prime}=\frac{1}{b}\sum_{i=1}^{b}z^{[m+i]}$ and $\upsilon^{\prime}_{b}=\frac{1}{b}\sum_{i=1}^{b}(z^{[m+i]}-\mu^{\prime}_{b})^{2}$.
@@
  RETURN: Monosementicity Scale of inputs $\phi({\mathbf{h}})$
@@
  Input: Values of $n$ neurons with batchsize $b$: $\mathbf{z}=\{\{z_{i}^{[j]}\}_{j=1}^{b}\}_{i=1}^{n}$.
@@
$\phi(h_{i}^{[m+j]})=\frac{(h_{i}^{[m+j]}-{\mu_{i}}_{m})^{2}}{{\upsilon_{i}}_{m}+\epsilon\sum_{k=1}^{n}{\upsilon_{k}}_{m}/n}\text{ for }i\in[1,n],j\in[1,b].$
     Select values with high $\phi$, adjust each of them according to MD. Replacing the original ones in $\mathbf{z}$ to form output $\mathbf{z}^{\prime}$.
@@
  RETURN: Adjusted layer of neurons $\mathbf{z}^{\prime}$
D.2. Implementation Details of MEmeL
Appendix E Experiment Setting
E.1. Configuration of Base Models
@@
For the precipitation task, we insert 6 MEemL lay- ers after the hidden states of each ConvGRU module. The training is conducted on the HKO-7 dataset (Shi et al., 2017) of radar images with a granularity of $120\times 120$, which are centered in Hong Kong. These radar images cover a timespan from 2009 to 2015 and are captured every 6 minutes. The original rainfall intensity is mapped to a scale of [0, 255], forming a regression task. The pretraining settings are the same with theirs, optimizing for 20k steps. Based on that, we conduct 2k more steps funetuning to obtain the final output.
E.2. Metrics
@@
  • The Accuracy: With the statistic of True Positive (TP), False Positive (FP), True Negative (TN), and False Negative (FN), Accuracy can be derived as $\frac{\text{TP+TN}}{\text{TP+TN+FN+FN}}$.
@@
E.3. Parameter Searching over Inhibition Level of Monosemanticity
As mentioned above, to ensure fairness, we apply the minimal scale of inhibition, wherein only the most monosemantic neuron is selected and deactivated for each sample. However, we can also adjust the inhibition scale in each step by introducing a factor $\Delta$. By unifying the two naive methods and our reverse deactivation, we obtain the following function:
$z^{\prime}=-\Delta\cdot z+(\bar{z}+\Delta\cdot z)_{ng}.$
Our reverse deactivation (RD) is a special case when $\Delta=1$, while the two naive methods can be obtained with $\Delta=0$ and -1. It can be easily proven that for any $\Delta>0$, the modification can deactivate neuron $z$ according to Lemma 3.3, and a larger $\Delta$ corresponds to a more aggressive deactivation. Therefore, we search for the number of neurons $n_{n}$ to deactivate in a single step and the scale of deactivation $\Delta$ for the best performance. We select the best model from the search space: $n_{n}\in\{1,5,10,50,100\}$ and $\Delta\in\{1,5,10,50,100\}$.
Table 5. Validation results on GLUE Test datasets. The settings are the same with Table. 1. All metrics are the larger the better with best results in bold font.
Model MNLI-(M/MM) QQP QNLI SST-2 CoLA STS-B MRPC RTE Average
MEmeL 84.8/83.9 71.7 90.9 93.6 54.5 86.6 87.6 66.4 80.0
MEmeL-T 84.9/83.7 71.5 90.5 93.9 52.1 85.9 88.7 68.1 79.9
Table 6. Validation results on ImageNet-1k dataset. See Table. 2 for detailed settings. The metric is the higher the better. The best results are indicated in bold font.
Model Swin-T Swin-S Swin-B
Size 28M 50M 88M
MEmeL 81.1 83.4 85.1
MEmeL-T 81.1 83.5 85.1
Table 7. Results on HKO-7 dataset. The settings are the same with Table. 2. The metrics are the smaller the better with the best results in bold fonts.
Model B-MAE B-MSE
MEmeL 1003.25 309.94
MEmeL-T 1002.46 309.52
Appendix F More Experiments
F.1. The Impact of Removing MEmeL During Test
Recall that in subsection 3.3, we emphasize the lightness of our MEmeL. It aims to induce models to reduce monosemanticity during training and can be removed during testing. Here, we conduct experiments to enable MEmeL during testing, which are expected to have similar results compared to the results displayed in the main experiments (see Section 4.2). These models are labeled with the postfix “-T” (e.g., MEmeL-T for MEmeL as the base model) while keeping other experimental settings unchanged.
The results are displayed in Table. 5, Table. 6, and Table. 7 for GLUE, ImageNet, and HKO-7 datasets, respectively. The best results are shown in bold fonts. The performance differences for the two settings are very close on all three tasks, indicating a stable performance when tested without the induction of MEmeL. Interestingly, on GLUE datasets, the performance when using MEmeL during testing is much more fluctuated. It would be valuable to study the dynamics of MEmeL in more depth as future work.
@@ Appendix G Discussions
G.1. Biological View of Neuron Connection
Monosemanticity and polysemanticity are inherent mappings between neurons, where decreasing monosemanticity encourages the construction of complex mappings instead of exclusive connections. With a similar functionality, in neuroscience, “synaptic” enables information transmission between neurons (Hasson et al., 2020; Thomas et al., 2017). In view of the evolution of the brain, the increase of neuron activity is positively correlated to the increase of the amount of synaptic (M and S, 2015). Besides, compared with chimpanzees, humans significantly have more synaptics for information transmission (A et al., 2012).
However, the connections of artificial neurons are fixed during design. In contrast, the connection relationships of biological neurons are observable, which can serve as a reference to study the monosemanticity of AI neurons in large-scale models.
G.2. Brain-inspired ANN design
@@
G.3. Difference from Traditional Normalizations
G.4. On the Selection of Benchmark and Experimental Settings
@@
However, validating the idea and performing pretraining is much more computationally expensive compared to finetuning, especially for large models. For instance, using the naive method to pretrain LLAMA2-7B based on our A800 is estimated to cost over \$$450k$ and over 5800 GPU-days (Touvron et al., 2023) with our preliminary test. Emergence occurs primarily near $10^{22}$ FLOPs and costs more than 10500 GPU-days (Wei et al., 2022). Conducting corresponding validation experiments is impossible for most researchers without the assistance of top companies.
Under such an arms race on GPUs, our experiments explore our resources to first validate the generality and effectiveness of our methods. Though finetuning may not fully utilize the capabilities of MEmeL, our main experiments demonstrate that finetuning with MEmeL is powerful enough to improve results. Additionally, on the precipitation forecasting task with a smaller scale, we conduct pretraining with MEmeL and observe a significant improvement compared to finetuning. The details are provided in subsection. 4.3. This finding supports our idea that compared to finetuning, pretraining can leverage the strengths of MEmeL better.
G.5. Potential Directions of Emergence Learning
Wei et al. (2022) raise the concept of emergence in AI models and point out that the explanation for emergence is still under investigation. They discuss the opinion that AI models can be seen as compressors and explain emergence as a result of the accumulation to compress world knowledge. In this aspect, decreasing monosemanticity encourages models to compress knowledge rather than storing knowledge that compromises generality.
G.6. Future Work of Partially Decreasing Monosemanticity
Increasing monosemanticity can be beneficial in some areas where precise memorization is necessary. It is important to study the impact of monosemanticity on different tasks and inhibit or increase it according to the analysis. A complex task could also be broken down into smaller steps, each with its own preference for monosemanticity.
For example, when asking a model to imitate the author of “Sapiens: A Brief History of Humankind” and write a preface from scratch, the model should accurately identify Harari as the author through a key-value style memorization and then reproduce the tone in the output with the necessary reasoning ability.